OpenStudy (anonymous):

describe the process of finding the equation of a line given the slope and a point and give an example

OpenStudy (triciaal):

slope,m point (x1, y1) slope of a line = change in y / corresponding change in x to find the slope you use any 2 points on the given line let the other point = (x2, y2) the x-intercept b occurs when y = 0 at the point (x1, y1) y1 = m x1 + b m = (y2 - y1)/ (x2 - x1) with this you can find b the intercept when x = 0 using this information write the equation of the line for any point on that line (x, y) y = m x + b the slope-intercept form or (y - y1) = m(x - x1) the point-slope form

OpenStudy (triciaal):

an example slope = 2 and passing through point (3, 7) (y -7) = 2(x - 3) point slope format y = 2 x+ ( -6 + 7 ) y = 2 x + 1 slope-intercept format

OpenStudy (triciaal):

I gave more than needed for this question. You were given the slope. do you have any questions about my response?

OpenStudy (anonymous):

no thank you that helps out alot!!! much appreciated!!